Determine the EAC given a 5% cost of capital with cash outflows starting at $30,000 in year 1, and grow by 3% per year for the next 8 years.
Answer #1

PV of costs =30000/(1+5%)+30000*(1+3%)/(1+5%)^2.................+30000*(1+3%)^7/(1+5%)^8
=30000/(1+5%)*(1+1.03/1.05+(1.03/1.05)^2+......(1.03/1.07)^7)=30000/1.05*(1-(1.03/1.05)^8)/(1-(1.03/1.05) =213900.22


EAC=NPV/((1-(1+r)^-n)/r) =213900.22/((1-(1+5%)^-8)/5%) =22400.22